2/3(7/8-4x)-5/8=3/8
Answer by rfer(16322)

2/3(7/8-4x)-5/8=3/8
14/24-8/3x-5/8=3/8
14/24-64/24x-15/24=9/24
-1/24-64/24x=9/24
-64/24x=10/24
x=(10/24)/(-64/24)
x=-240/1536
x=-5/32